A candy store owner wishes to mix 30 pounds of candy worth $6.00 per pound with nuts worth $3.00 per pound to get a mixture worth $5.00 per pound. How many pounds of nuts must he have?
Answer by ptaylor(2198)

Let x=amount of nuts needed
Now we know that the value of the nuts(3x) and candy(6*30) before they are mixed has to equal the value of the nuts and candy after they are mixed(5*(30+x)), ok? ($ are understood) So, our equation to solve is:
3x+6*30=5(30+x) get rid of parens and simplify
3x+180=150+5x subtract 150 and also 3x from each side
3x-3x+180-150=150-150+5x-3x collect like terms
30=2x divide each side by 2
x=15 lb-------------amount of nuts needed
CK
3*15+6*30=5*45
45+180=225
225=225
Hope this helps---ptaylor